Validate that air has molecular weight of 28.97kg/(k mol).
1
Expert's answer
2020-11-23T05:37:18-0500

Air composition: 78 % N2, 21 % O2 and 1 % Ar.

M(O2) = 32 g/mol

M(N2) = 28 g/mol

M(Ar) = 40 g/mol

"M(Air) = 28 \times 0.78 + 32 \times 0.21 + 40 \times 0.01 = 21.84 + 6.72 + 0.4 = 28.96 \;g/mol \;or\; kg/(k\;mol)"